I am hung up on a word problem. Can anyone help? It is:
A rectangle with length 2x-1 inches has an area of 2x^2 + 5x - 3 square inches. Write a binomial that represents its width.
Thank you for any help you can give me.

Answer by venugopalramana(3286)

A rectangle with length 2x-1 inches has an area of 2x^2 + 5x - 3 square inches. Write a binomial that represents its width.
Thank you for any help you can give me.
AREA=LENGTH*WIDTH
(2X-1)*WIDTH=2X^2+5X-3=2X^2+6X-X-3=2X(X+3)-1(X+3)=(2X-1)(X+3)
WIDTH=(2X-1)(X+3)/(2X-1)=X+3